Suppose you enter into a 4.0 month forward contract on one ounce of silver when

the spot price of silver is $8.90 per ounce and the risk-free interest rate is 5.25

percent continuously compounded. What is the forward price?
Answer #1

Spot price of silver per ounce = $8.90

Continuously compounded rate = 5.25%

Time = 4 months = 4/12 = 0.33 years

Forward Price = Spot Price* e *T

= 8.90 * p5.25 +0.38 = $9.06